Please help me with this word problem on finding the average speed...
Julio left the hospital and drove toward his friend's house. One hour later Chelsea left driving 10 km/h faster in an effort to catch up to him. After four hours Chelsea finally caught up. Find Julio's average speed.

Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

let s = J's speed
then
(s+10) = C's speed
:
Julio left the hospital and drove toward his friend's house.
One hour later Chelsea left driving 10 km/h faster in an effort to catch up to him.
After four hours Chelsea finally caught up.
Find Julio's average speed.
:
We know that from the information given, that that:
J's travel time = 5 hrs (1 hr more than C's)
C's travel time = 4 hrs
:
We know that when C catches J, they will have traveled the same distance.
Write a distance equation: dist = time * speed
J's dist = C's dist
5s = 4(s+10)
5s = 4s + 40
5s - 4s = 40
s = 40 mph is J's speed
:
:
:
Confirm this solution, find the actual dist of each
5(40) = 200 mi
4(40+10) = 200 mi

Answer by ikleyn(52798)

.
There is a special lesson on catching up problems in this site
    - Travel and Distance problems for two bodies moving in the same direction (catching up)

You can find different types of solved catching up problems there.